OpenStudy (anonymous):

(square root of a+1 plus the square root of a-1) (SR of a+1 minus the SR of a-1)

OpenStudy (anonymous):

You have the difference of square, which in the general form is (a^2)-(b^2)=(a+b)(a-b). In this scenario, your a=sqrt(a+1) and b=sqrt(a-1), which means this is just 2.